# Euler-Heisenberg black hole surrounded by quintessence in the background of perfect fluid dark matter: Thermodynamics, Shadows and Quasinormal modes

### v1

@@ Abstract
Department of Physics, University of Hradec Králové, Faculty of Science, Rokitanského 62, Hradec Kralove, Královéhradecký , 500 03, Czech Republic
@@ 1 Introduction
Indirect but reliable evidence, such as the rotation curves of spiral galaxies [21, 22, 23], the dynamical stability of disk galaxies [25, 24], the formation of large-scale structure in the Universe [26], gravitational lensing observations of galaxies and galaxy clusters [27, 28], the cosmic microwave background [29, 30], and baryon acoustic oscillations [31], points out that a considerable part of the Universe is made up of dark matter, distinct from ordinary observable matter and dark energy [32, 33]. Therefore, especially in the last two decades, scientists have proposed many theories to model dark matter, which goes or not beyond the Standard Model [34, 35, 36, 37, 38, 39, 40, 44, 41, 42, 43, 45]. So far, cold, warm and scalar field dark matter models have been considered as suitable dark matter models [46, 47, 48]. In addition to these, a new model, Perfect Fluid Dark Matter (PFDM) [49], has been added to the list of suitable models because of its success in describing the asymptotically flat rotation curves of observed spiral galaxies [50, 51]. Although in some aspects the real dark matter could not fulfill the perfect fluid condition, still it has been widely assumed as a consistent approach, and it has been used widely especially in recent years [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
@@ 2 A brief review
$\mathcal{L}\left(F,G\right)=-F+\frac{a}{2}F^{2}+\frac{7a}{8}G^{2}.$ (2)
Here, $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu},$ $G=\frac{1}{4}F_{\mu\nu}^{\text{ \ \ }\ast}F^{\mu\nu}$ and $a$ is the Euler-Heisenberg parameter that regulates the strength of the NLED contribution. The equations of motion derived from this action are more conveniently expressed using the Legendre dual description of NLED, which was introduced by Plebański in [92]. This approach involves the introduction of a tensor $P_{\mu\nu}$,
@@
To the first order in the constant $a$, the Hamiltonian of the Euler-Heisenberg theory reads
@@
$\mathbf{S=}\frac{1}{4\pi}\int d^{4}x\sqrt{-g}\left[\frac{1}{4}R-\left(\mathcal{L}\left(F,G\right)+4\pi\mathcal{L}_{DM}-\mathcal{L}_{qui}\right)\right].$ (8)
Here, the dark matter contribution to the action is represented by the term $\mathcal{L}_{DM}$, which is the Lagrangian density for dark matter. The quintessence term is denoted by $\mathcal{L}_{qui}$ and given in the form of [94]:
$\mathcal{L}_{qui}=-\frac{1}{2}\left(\nabla\phi\right)^{2}-V\left(\phi\right),$ (9)
@@
$\nabla_{\mu}P^{\mu\nu}=0,\text{ \ \ \ \ }R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=8\pi(\bar{T}_{\mu\nu}+4\pi T_{\mu\nu}(DM)-T_{\mu\nu}(qui)),$ (10)
@@
$T_{\mu\nu}(qui)$ $=$ $\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\left[\left(\nabla\phi\right)^{2}+2V\left(\phi\right)\right],$ (11)
$T_{\mu\nu}(DM)$ $=$ $\frac{2\delta\left(\sqrt{-g}\mathcal{L}_{DM}\right)}{\delta g^{\mu\nu}}.$ (12)
@@
In the simplest scenario, the energy density to pressure ratio is assumed to be constant, $\frac{\rho}{p}=\delta-1$, [52]. Next, we determine the solution to the field equations for a static, spherically symmetric metric that takes the following form:
@@
$f(r)=1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}-\frac{aq^{4}}{20r^{6}}-\frac{\sigma}{r^{3\omega_{q}+1}}+\frac{\alpha}{r}\log\left(\frac{r}{\alpha}\right),$ (19)
where $\sigma$, $\omega_{q}$ and $\alpha$ correspond to the quintessence matter normalization factor, quintessence state parameter and the PFDM intensity, respectively. Before examining the thermal quantities, we would like to show qualitatively in graphs how quintessence matter and PFDM change the metric function. To do this, we first plot the metric function versus radius with three different quintessence state parameter values in Fig. 1.
@@ 3 Thermodynamics
$M_{H}=\frac{r_{H}}{20}\left(1+\frac{q^{2}}{r_{H}^{2}}-\frac{aq^{4}}{20r_{H}^{6}}-\frac{\sigma}{r_{H}^{1+3\omega_{q}}}+\frac{\alpha}{r_{H}}\log\left(\frac{r_{H}}{\alpha}\right)\right).$ (20)
@@ 4 Black hole Shadow
The black hole shadow holds immense significance in astrophysics as it offers a chance to validate Einstein’s theory of general relativity. The shadow’s dimensions and shape are governed by the curvature of spacetime surrounding the black hole, a key prediction of this theory. Through careful observation and analysis of the shadow, researchers can examine the theory’s predictions, seeking any anomalies that might enhance our comprehension of the fundamental physical laws governing our Universe. To study the shadow, our first step is to explore variations in the photon sphere. Utilizing Eq. (19) and following the conventional methodology outlined in [95], we can numerically solve the equation to ascertain the precise location of the photon sphere. The equations governing the movement of photons in spacetime, dictated by the metric (19), can be derived from the Lagrangian formulation,
@@
$\dot{r}^{2}+\mathcal{V}_{eff}\left(r\right)=0,$ (44)
where $\mathcal{V}_{eff}\left(r\right)$ is the effective potential for radial photon’s motion defined by,
$\mathcal{V}_{eff}\left(r\right)=\left(\frac{\mathcal{K}+L^{2}}{r^{2}}\right)\left(1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}-\frac{aq^{4}}{20r^{6}}-\frac{\sigma}{r^{3\omega_{q}+1}}+\frac{\alpha}{r}\log\left(\frac{r}{\alpha}\right)\right)-E^{2}.$ (45)
@@
$\mathcal{V}_{eff}\left(r_{ph}\right)=0,\quad\text{or}\quad\mathcal{R}\left(r_{ph}\right)=0,$ (46)
@@
$\left.\frac{d}{dr}\mathcal{V}_{eff}\left(r\right)\right|_{r=r_{ph}}=0,\quad\text{or}\quad\left.\frac{d}{dr}\mathcal{R}\left(r\right)\right|_{r=r_{ph}}=0.$ (47)
Here, the photon sphere radius, $r_{ph}$, refers to the specific value of $r$ at which the effective potential $\mathcal{V}_{eff}\left(r_{ph}\right)$ reaches its maximum value ($\left.\frac{d^{2}}{dr^{2}}\mathcal{V}_{eff}\left(r\right)\right|_{r=r_{ph}}<0$). Based on the conditions introduced in Eq. (47), the photon sphere radius can be determined by solving the following equation:
$\frac{2aq^{4}}{5r_{ph}^{6}}+\frac{\alpha}{r_{ph}}+\frac{6M}{r_{ph}}+\frac{3\sigma(\omega_{q}+1)}{r_{ph}^{1+3\omega_{q}}}-\frac{4q^{2}}{r_{ph}^{2}}-\frac{3\alpha}{r_{ph}}\log\left(\frac{r_{ph}}{\alpha}\right)-2=0.$ (48)
@@
$\xi^{2}+\eta=\frac{r_{ph}^{2}}{1-\frac{2M}{r_{ph}}+\frac{q^{2}}{r_{ph}^{2}}-\frac{aq^{4}}{20r_{ph}^{6}}-\frac{\sigma}{r_{ph}^{3\omega_{q}+1}}+\frac{\alpha}{r_{ph}}\log\left(\frac{r_{ph}}{\alpha}\right)}.$ (50)
@@
Table 5 shows how $\eta+\xi^{2}$ varies with respect to the dark matter intensity and the quintessence state parameters. From this table, we can see that for a constant value of $\omega_{q}$, the value of $\eta+\xi^{2}$ decreases as $\alpha$ increases. On the other hand, if we fix the value of $\alpha$, the value of $\eta+\xi^{2}$ increases as the quintessence parameter increases. Moreover, in Table 6, we present the computed values of the photon radius $r_{ph}$ and $\eta+\xi^{2}$ for different values of $\sigma$ and $\omega_{q}$. We observe that the impact of the quintessence normalization factor does not mimic mimics the effect of the dark matter intensity parameter. More precisely, as $\sigma$ increases, the photo radius and the $\eta+\xi^{2}$ increases.
@@
$\frac{d\varphi}{dr}$ $=$ $\frac{\xi}{\sin^{2}\theta_{0}\sqrt{r_{0}^{4}-r_{0}^{2}f\left(r_{0}\right)\left(\xi^{2}+\eta\right)}},$ (53)
$\frac{d\theta}{dr}$ $=$ $\frac{\sqrt{\eta-\xi^{2}\cot^{2}\theta_{0}}}{\sqrt{r_{0}^{4}-r_{0}^{2}f\left(r_{0}\right)\left(\xi^{2}+\eta\right)}}.$ (54)
@@
$X$ $=$ $-\frac{\zeta}{\sin\theta_{0}},$ (55)
$Y$ $=$ $\sqrt{\eta-\zeta^{2}\cot^{2}\theta_{0}}.$ (56)
Then, in the equatorial plane $\theta_{0}=\pi/2$, the celestial coordinates can be written as
$X^{2}+Y^{2}=\eta+\zeta^{2}=R_{S}^{2}.$ (57)
Here, $R_{S}$ is the radius of the black hole shadow. We now use the equations (51) and (52) to plot the contour of the Euler-Heisenberg black hole shadow at the equatorial plane for different values of the quintessence parameter and dark matter in Fig. 12 and Fig. 13.
@@
Fig. 13 reveals the impact of the quintessence normalization factor. Unlike the dark matter intensity effect, the shadow radius increases as $\sigma$ increases. Moreover, we observe that this increase is greater at smaller $\omega_{q}$ values. To consolidate our results, we present Fig. 14 and Fig. 15.
@@
Figure 14: Variation of the radius of the black hole shadow $R_{s}$ vs. $\omega_{q}$ for $M=1$, $q=0.5$, $a=0.1$.
@@
Figure 15: Variation of the radius of the black hole shadow $R_{s}$ vs. $\sigma$ /or $\alpha$ for $M=1$, $q=0.5$, $a=0.1$.
@@ 5 Quasinormal modes: Wentzel-Kramers-Brillouin approximation
Over the past few decades, the QNMs have been investigated using various methods. Each method has its strengths and weaknesses. Developing more efficient techniques to study QNMs is an active area of ongoing research. In the present study, we employ the semi-analytical Wentzel-Kramers-Brillouin (WKB) approximation method. This technique involves matching the WKB expansion of the modes at the event horizon and spatial infinity with the Taylor expansion of the effective potential near the peak of the potential barrier. The WKB method was initially proposed in [97, 98, 99, 100] and subsequently extended to higher-order approximations, including the 6th order [101, 102, 103] and 13th order [104]. It is noteworthy that increasing the order of the WKB approximation does not always result in a better approximation for the frequency. Consequently, we have chosen to utilize the 6th order expansion for our study, which is represented by the following equation,
@@
Fig. 18 illustrates how dark matter changes the real part ($Re\omega$) and the negative imaginary part ($-Im\omega$) of the frequency. We observe that the real part becomes larger as the value of $n$ increases. We note that the plots for the negative imaginary part show a similar trend.

### v2

@@ Abstract
Department of Physics, Faculty of Science, University of Hradec Kralove,
Rokitanskeho 62/26, Hradec Kralove, 500 03, Czech Republic.
@@ 1 Introduction
Indirect but reliable evidence, such as the rotation curves of spiral galaxies [21, 22, 23], the dynamical stability of disk galaxies [25, 24], the formation of large-scale structure in the Universe [26], gravitational lensing observations of galaxies and galaxy clusters [27, 28], the cosmic microwave background [29, 30], and baryon acoustic oscillations [31], points out that a considerable part of the Universe is made up of dark matter, distinct from ordinary observable matter and dark energy [32, 33]. Therefore, especially in the last two decades, scientists have proposed many theories to model dark matter, which goes or not beyond the Standard Model [34, 35, 36, 37, 38, 39, 40, 44, 41, 42, 43, 45]. So far, cold, warm and scalar field dark matter models have been considered as suitable dark matter models [46, 47, 48]. In addition to these, a new model, perfect fluid dark matter (PFDM) [49, 50, 51], has been added to the list of suitable models because of its success in describing the asymptotically flat rotation curves of observed spiral galaxies [52, 53]. Although in some aspects the real dark matter could not fulfill the perfect fluid condition, still it has been widely assumed as a consistent approach, and it has been used widely especially in recent years [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 68, 66, 67, 69, 70, 71, 72, 73].
@@ 2 A brief review
$\mathcal{L}\left(F,G\right)=-F+\frac{a}{2}F^{2}+\frac{7a}{8}G^{2},$ (2)
while
$F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\qquad\text{and}\qquad G=\frac{1}{4}F_{\mu\nu}^{\text{ \ \ }\ast}F^{\mu\nu}.$ (3)
Here, $a$ is the Euler-Heisenberg parameter that regulates the strength of the NLED contribution and it is defined with the fine structure constant, $\alpha_{fs}$, and the electron mass, $m$, in the natural units ($\hbar=c=1$) as follows:
$a=\frac{8\alpha_{fs}}{45m^{4}}.$ (4)
The equations of motion derived from this action are more conveniently expressed using the Legendre dual description of NLED, which was introduced by Plebański in [95]. This approach involves the introduction of a tensor $P_{\mu\nu}$,
@@
In circumstances where the electric fields are relatively weak in comparison to the critical field strength $E_{c}$,
$\frac{\alpha_{fs}}{E_{c}}<<1,$ (7)
a perturbative expansion of the Euler-Heisenberg Lagrangian becomes a valid approach. Since $a$ is small, higher-order corrections in $a$ are negligible compared to the first-order term. It is important to note that this assumption is always valid unless the system is subjected to extreme conditions, such as the presence of very strong fields or highly charged black holes. Based on this premise, we express the Hamiltonian in the first order in $a$ as follows:
@@
$\mathbf{S=}\frac{1}{4\pi}\int d^{4}x\sqrt{-g}\left[\frac{1}{4}R-\left(\mathcal{L}\left(F,G\right)+4\pi{\mathcal{L}_{\rm DM}}-{\mathcal{L}_{\rm qui}}\right)\right].$ (11)
Here, the dark matter contribution to the action is represented by the term $\mathcal{L}_{\rm DM}$, which is the Lagrangian density for dark matter. The quintessence term is denoted by $\mathcal{L}_{\rm qui}$ and given in the form of [97]:
${\mathcal{L}_{\rm qui}}=-\frac{1}{2}\left(\nabla\phi\right)^{2}-V\left(\phi\right),$ (12)
@@
$\nabla_{\mu}P^{\mu\nu}=0,\text{ \ \ \ \ }R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=8\pi(\bar{T}_{\mu\nu}+4\pi T^{(\rm DM)}_{\mu\nu}-T^{(\rm qui)}_{\mu\nu}),$ (13)
@@
$T^{(\rm qui)}_{\mu\nu}$ $=$ $\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\left[\left(\nabla\phi\right)^{2}+2V\left(\phi\right)\right],$ (14)
$T^{(\rm DM)}_{\mu\nu}$ $=$ $\frac{2\delta\left(\sqrt{-g}\mathcal{L}_{\rm DM}\right)}{\delta g^{\mu\nu}}.$ (15)
@@
In the simplest scenario, the energy density to pressure ratio is assumed to be constant, $\frac{\rho}{p}=\delta-1$, [54]. Here, we posit that the dark energy of the universe is predominantly constituted by the potential of a scalar field in the form of [98]
$V\left(\phi\right)\simeq e^{-\sqrt{\frac{3}{2}\left(1+\omega_{q}\right)}\phi},$ (17)
which leads to a special solution such that the equation of state reads $P=\omega_{q}\rho.$ We now determine the solution to the field equations for a static, spherically symmetric metric that takes the following form:
@@
$f(r)=1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}-\frac{aq^{4}}{20r^{6}}-\frac{\sigma}{r^{3\omega_{q}+1}}+\frac{\alpha}{r}\log\left(\frac{r}{{\lvert\alpha\rvert}}\right),$ (23)
where $\sigma$, $\omega_{q}$ , $\alpha$ and $q$ correspond to the quintessence matter normalization factor, quintessence state parameter, PFDM intensity and the electric charge of black hole, respectively. Before examining the thermal quantities, we would like to show qualitatively in graphs how quintessence matter and PFDM change the metric function. To do this, we first plot the metric function versus radius with three different quintessence state parameter values in Fig. 1.
@@ 3 Thermodynamics
$M_{H}={\frac{r_{H}}{2}}\left(1+\frac{q^{2}}{r_{H}^{2}}-\frac{aq^{4}}{2r_{H}^{6}}-\frac{\sigma}{r_{H}^{1+3\omega_{q}}}+\frac{\alpha}{r_{H}}\log\left(\frac{r_{H}}{{\lvert\alpha\rvert}}\right)\right).$ (24)
@@ 4 Black hole Shadow
The black hole shadow holds immense significance in astrophysics as it offers a chance to validate Einstein’s theory of general relativity. The shadow’s dimensions and shape are governed by the curvature of spacetime surrounding the black hole, a key prediction of this theory. Through careful observation and analysis of the shadow, researchers can examine the theory’s predictions, seeking any anomalies that might enhance our comprehension of the fundamental physical laws governing our Universe [99, 100, 101, 102, 103, 104, 105, 106, 107]. To study the shadow, our first step is to explore variations in the photon sphere. Utilizing Eq. (23) and following the conventional methodology outlined in [108], we can numerically solve the equation to ascertain the precise location of the photon sphere. The equations governing the movement of photons in spacetime, dictated by the metric (23), can be derived from the Lagrangian formulation,
@@
$\dot{r}^{2}+\mathcal{V}_{{\rm eff}}\left(r\right)=0,$ (48)
where $\mathcal{V}_{{\rm eff}}\left(r\right)$ is the effective potential for radial photon’s motion defined by,
$\mathcal{V}_{{\rm eff}}\left(r\right)=\left(\frac{\mathcal{K}+L^{2}}{r^{2}}\right)\left(1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}-\frac{aq^{4}}{20r^{6}}-\frac{\sigma}{r^{3\omega_{q}+1}}+\frac{\alpha}{r}\log\left(\frac{r}{{\lvert\alpha\rvert}}\right)\right)-E^{2}.$ (49)
@@
$\mathcal{V}_{{\rm eff}}\left(r_{{\rm ph}}\right)=0,\quad\text{or}\quad\mathcal{R}\left(r_{{\rm ph}}\right)=0,$ (50)
@@
$\left.\frac{d}{dr}\mathcal{V}_{{\rm eff}}\left(r\right)\right|_{r=r_{{\rm ph}}}=0,\quad\text{or}\quad\left.\frac{d}{dr}\mathcal{R}\left(r\right)\right|_{r=r_{{\rm ph}}}=0.$ (51)
Here, the photon sphere radius, $r_{{\rm ph}}$, refers to the specific value of $r$ at which the effective potential $\mathcal{V}_{{\rm eff}}\left(r_{{\rm ph}}\right)$ reaches its maximum value ($\left.\frac{d^{2}}{dr^{2}}\mathcal{V}_{{\rm eff}}\left(r\right)\right|_{r=r_{{\rm ph}}}<0$). Based on the conditions introduced in Eq. (51), the photon sphere radius can be determined by solving the following equation:
$\frac{2aq^{4}}{5r_{{\rm ph}}^{6}}+\frac{\alpha}{r_{{\rm ph}}}+\frac{6M}{r_{{\rm ph}}}+\frac{3\sigma(\omega_{q}+1)}{r_{{\rm ph}}^{1+3\omega_{q}}}-\frac{4q^{2}}{r_{{\rm ph}}^{2}}-\frac{3\alpha}{r_{{\rm ph}}}\log\left(\frac{r_{{\rm ph}}}{{\lvert\alpha\rvert}}\right)-2=0.$ (52)
@@
$\xi^{2}+\eta=\frac{r_{{\rm ph}}^{2}}{1-\frac{2M}{r_{{\rm ph}}}+\frac{q^{2}}{r_{{\rm ph}}^{2}}-\frac{aq^{4}}{20r_{{\rm ph}}^{6}}-\frac{\sigma}{r_{{\rm ph}}^{3\omega_{q}+1}}+\frac{\alpha}{r_{{\rm ph}}}\log\left(\frac{r_{{\rm ph}}}{{\lvert\alpha\rvert}}\right)}.$ (54)
@@
Table 5 shows how $\eta+\xi^{2}$ varies with respect to the dark matter intensity and the quintessence state parameters. From this table, we can see that for a constant value of $\omega_{q}$, the value of $\eta+\xi^{2}$ decreases as $\alpha$ increases. On the other hand, if we fix the value of $\alpha$, the value of $\eta+\xi^{2}$ increases as the quintessence parameter increases. Moreover, in Table 6, we present the computed values of the photon radius $r_{{\rm ph}}$ and $\eta+\xi^{2}$ for different values of $\sigma$ and $\omega_{q}$. We observe that the impact of the quintessence normalization factor does not mimic the effect of the dark matter intensity parameter. More precisely, as $\sigma$ increases, the photon radius and the $\eta+\xi^{2}$ increases.
@@
$\frac{d\varphi}{dr}$ $=$ $\frac{\xi}{\sin^{2}\theta_{o}\sqrt{r_{o}^{4}-r_{o}^{2}f\left(r_{o}\right)\left(\xi^{2}+\eta\right)}},$ (57)
$\frac{d\theta}{dr}$ $=$ $\frac{\sqrt{\eta-\xi^{2}\cot^{2}\theta_{o}}}{\sqrt{r_{o}^{4}-r_{o}^{2}f\left(r_{o}\right)\left(\xi^{2}+\eta\right)}}.$ (58)
@@
$X$ $=$ $-\frac{\zeta}{\sin\theta_{o}},$ (59)
$Y$ $=$ $\sqrt{\eta-\zeta^{2}\cot^{2}\theta_{o}}.$ (60)
Then, in the equatorial plane $\theta_{o}=\pi/2$, the celestial coordinates can be written as
$X^{2}+Y^{2}=\eta+\zeta^{2}=R_{{\rm S}}^{2}.$ (61)
Here, $R_{{\rm S}}$ is the radius of the black hole shadow. We now use the equations (55) and (56) to plot the contour of the Euler-Heisenberg black hole shadow at the equatorial plane for different values of the quintessence parameter and dark matter in Fig. 12 and Fig. 13.
@@
Fig. 13 reveals the impact of the quintessence normalization factor. Unlike the dark matter intensity effect, the shadow radius increases as $\sigma$ increases. Moreover, we observe that this increase is greater at smaller $\omega_{q}$ values. To consolidate our results, we provide additional visualization. For example in Fig. 14, we present a comparison of the changes in black hole shadow radius.
@@
Figure 14: Variation of the radius of the black hole shadow $R_{\rm S}$ versus $\omega_{q}$ for $M=1$, $q=0.5$, $a=0.1$.
For a fixed value of sigma, we observe that as the quintessence state parameter increases the shadow radius decreases at a slower rate at larger alpha values. In contrast to this, for a fixed value of $\alpha$, we see that as $\omega_{q}$ increases, the shadow radius decreases at a greater rate at larger $\sigma$ values. In Fig. 15, we demonstrate a comparison of the alteration in shadow radius versus dark matter and energy parameters.
@@
Figure 15: Variation of the radius of the black hole shadow $R_{\rm S}$ versus $\sigma$ /or $\alpha$ for $M=1$, $q=0.5$, $a=0.1$.
We observe that for a fixed value of $\sigma$, as $\alpha$ increases, the radius decreases at a greater rate for smaller values of the quintessence state parameter. In contrast, for a fixed value of $\alpha$, as $\sigma$ decreases, the radius increases at a greater rate for greater values of the quintessence state parameter.
@@ 5 Quasinormal modes: Wentzel-Kramers-Brillouin approximation
Over the past few decades, the QNMs have been investigated using various methods [110]. Each method has its strengths and weaknesses. Developing more efficient techniques to study QNMs is an active area of ongoing research. In the present study, we employ the semi-analytical Wentzel-Kramers-Brillouin (WKB) approximation method. This technique involves matching the WKB expansion of the modes at the event horizon and spatial infinity with the Taylor expansion of the effective potential near the peak of the potential barrier. The WKB method was initially proposed in [111, 112, 113, 114] and subsequently extended to higher-order approximations, including the 6th order [115, 116, 117] and 13th order [118, 119]. It is noteworthy that increasing the order of the WKB approximation does not always result in a better approximation for the frequency. Consequently, we have chosen to utilize the 6th order expansion for our study, which is represented by the following equation,
@@
[…]